7x2-x+3k=0 Using quadradic equation, solve for k.
Answer by ankor@dixie-net.com(22740)

7x^2 - x + 3k = 0
:
Find the factors that satisfy the coefficients of x^2 and x
(7x + 6)(x - 1) = 0
Foil
7x^2 - 7x + 6x - 6 = 0
therefore
3k = -6
k = -6/3
k = -2